OpenStudy (anonymous):

Part 1: Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. (3 points) 5x + 3y = –7 4x – 2y = –10 Part 2: Show your work and provide the solution to the system.

OpenStudy (anonymous):

i would use the elimination method to about too many fraction multiply the first equation by 2 and the 2nd equation by 3 10x+6y = -14 12x-6y = -30 add the two equations to get 22x = -44 x=-2 replace x into the equation -10 +3y =-7 3y=3 y=1 x=-2

OpenStudy (anonymous):

Wow, thank you SO much! You're a life saver!

OpenStudy (anonymous):

I would give you 100 medals right now if I could!

OpenStudy (anonymous):

and im not sure what i intended to say but not the word about haha

OpenStudy (anonymous):

avoid! thats the word

OpenStudy (anonymous):

lol, alrightt. Btw, I have another question. Do you think you can answer it too? :/

OpenStudy (anonymous):

well whats the question haha but i will try

OpenStudy (anonymous):

Alright sweet! Let me post it now, you should see it with all the other questions. I'll give you another medal if I can by the way. lol